OpenStudy (anonymous):

hey guys was wondering if i could get help for finals i have two questions i need help with. 1.Stephanie lost 45 pounds in 3 months. the first month she lost ten pounds more than the third month.The second monnth she lost five pounds more than the third month.How many pounds did she lose the second month?

OpenStudy (anonymous):

im really bad at word problems i dont know how to set it up

OpenStudy (nikato):

Ok A=1sr B=2nd C=3rd month

OpenStudy (nikato):

A+B+C=45 A=10+C B=C-5 And ur solvingfor B

OpenStudy (anonymous):

whats c how do i find c?

OpenStudy (nikato):

Plug the value of A and B from the 2nd and 3rd equation into the first to find C

OpenStudy (anonymous):

oh i get it hang on

OpenStudy (anonymous):

so would my ending set up be 3c+5=45

OpenStudy (nikato):

10+c+c+5-c=45 c+15=45

OpenStudy (anonymous):

oh for the last c i put plus thats why

OpenStudy (nikato):

Wait, hold up

OpenStudy (nikato):

It's 10+c+c-5+c=45 3c+5=45 Ur correct

OpenStudy (anonymous):

i mean 40

OpenStudy (anonymous):

so c =8

OpenStudy (nikato):

Should be 3c=40 So c=40/3

OpenStudy (nikato):

Wait, how did u get 8

OpenStudy (anonymous):

thanks can you help me quickly with one more its the last question.

OpenStudy (anonymous):

are you sure thats the anwser cause my teacher got 10

OpenStudy (nikato):

Oh wait, misread the 3rd equation should be B=C+5

OpenStudy (nikato):

So 10+C+C+5+C=45 15+3C=45 3C=30 C=10